Ceramics Etc. is a manufacturer of large flower pots for urban settings. The company has these​ standards:

i Standard Price and Volume Direct materials (resin) 12 pounds per pot at a cost of $3.00 per pound 2.0 hours at a cost of $15.00 per hour Direct labor Standard variable manufacturing overhead rate $3.00 per direct labor hour Budgeted fixed manufacturing overhead $27,400 $8.00 per direct labor hour (DLH) Standard fixed MOH rate - X Actual Results Ceramics Etc. allocated fixed manufacturing overhead to production based on standard direct labor hours. Last month, the company reported the following actual results for the production of 1,900 flower pots Purchased 23,980 pounds at a cost of Direct materials $3.30 per pound; Used 23,180 pounds to produce 1,900 pots Worked 2.2 hours per flower pot (4,180 Direct labor total DLH) at a cost of $14.00 per hour $3.70 per direct labor hour for total Actual variable manufacturing overhead actual variable manufacturing overhead of $15,466 $27,100 Actual fixed manufacturing overhead Standard fixed manufacturing overhead $30,400 allocated based on actual production X
Requirement 1. Compute the variable manufacturing overhead variances. What do each of these variances tell management? (Enter the variances as positive numbers. Enter the currency amounts in the formulas to the nearest cent, then round the final variance amounts favorable (F) or unfavorable (U)) the nearest whole dollar. Label the variance as Begin by computing the variable manufacturing overhead rate variance. First determine the formula for the rate variance, then compute the rate variance for variable manufacturing overhead. Variable overhead rate variance X i Now compute the variable manufacturing overhead efficiency variance. First determine the formula for the efficiency variance, then compute the efficiency variance for variable manufacturing overhead. Variable overhead = efficiency variance X I What do each of these variances tell management? than expected. The The Vwere variable manufacturing overhead (MOH) rate variance tells managers that than variable MOH efficiency variance tells managers that actual were Requirement 2. Compute the fixed manufacturing overhead variances. What do each of these variances tell management? (Abbreviations used: MOH overhead. Enter the variances as positive numbers. Label the variances as favorable (F) or unfavorable (U).) manufacturing Begin by computing the fixed manufacturing overhead budget variance. First determine the formula for the budget variance, then compute the budget variance for fixed manufacturing overhead. Fixed MOH budget variance Now compute the fixed manufacturing overhead volume variance. First determine the formula for the volume variance, then compute the volume variance for fixed manufacturing overhead. Fixed MOH volume variance What do each of these variances tell management? fixed overhead budget variance tells managers that fixed The The overhead volume variance tells managers that
Answer #1

variable overhead rate variance

formula:

( actual rate - standard rate ) * actual hour

actual rate =3.7

standard rate = 3

actual hour = 4180

(3.7 - 3 ) * 4180 = 2926

variable rate variance = 2926 it is unfavorable ( actual rate > standard rate )

.

.

variable efficiency variance

formula:

( actual hour - standard hour ) * standard price

actual hour = 4180

standard hour = 3800 ( convert in to flexible = 2 standard hour * 1900 pots)

standard rate = 3

( 4180 - 3800 ) * 3= 1140

variable efficiency variance = 1140 it is unfavorable ( AH > SH )

3) FIXED OVERHEAD BUDGETED VARIANCE

actual fixed overhead - applied overhead

actual fixed overhead = 27100

applied fixed overhead = 30400 ( standard cost per hour fixed cost = 8 , and 2 hour needed one product so total cost per unit = 16 ( 8 * 2 ) , applied actual production 16 * 1900 = 30400 )

30400 - 27100 = 3300

fixed overhead budgeted variance = 3300 it is favorable ( A FOH < B FOH )

.

.

.

FIXED OVERHEAD VOLUME VARIANCE

BUDGETED FIXED OVERHEAD - APPLIED FIXED OVERHEAD

budgeted fixed overhead = 27400

applied fixed overhead = 30400

30400 - 27400 = 3000

fixed overhead volume variance = 3000 it is favorable